# Heterogeneous Treatment Effect in Time-to-Event Outcomes: Harnessing Censored Data with Recursively Imputed Trees

### v2

@@ 1 Introduction
In this paper, we introduce MISTR (Multiple Imputations for Survival Treatment Response), a new estimator for HTE estimation right-censored survival data. Inspired by Recursively Imputed Survival Trees (RIST) (Zhu and Kosorok, 2012), MISTR eliminates the need to estimate the censoring distribution. This in turn allows the flexible use of methods which otherwise would not be applicable to right-censored data. We build upon this flexibility and show how one can create methods which are more robust, more accurate, and applicable to a broader settings compared with existing methods. Figure 1 presents our approach schematically, where censored observations are efficiently imputed multiple times, treatment effects are estimated for each imputed dataset, and the resulting estimators are then combined.
@@
Cui et al. (2023) simulated settings 1–4 $\checkmark$ $\checkmark$ + $\checkmark$ +
Censoring probability reaches low values 6 $\times$ $\times$ $\checkmark$ +
Censoring depends on unobserved covariates 7 $\checkmark$ $\checkmark$ $\checkmark$ +
High censoring rate 8–9 $\checkmark$ $\checkmark$ $\checkmark$ +
Unobserved confounding with IV adjustment 201–204 $\checkmark$ $\times$ $\checkmark$ +
@@ 2 Related Work
To estimate causal effects in the presence of confounding, IV methods can be employedAngrist et al. (1996); Abadie (2003); Hernán and Robins (2006). An IV variable is associated with treatment assignment, affects the outcome only through the treatment, and is independent of unmeasured confounders. In survival data, IV methods have been widely studied. For example, Tchetgen Tchetgen et al. (2015) proposed a regression-based IV approach using additive hazard models, while Sørensen et al. (2019) and MacKenzie et al. (2014) applied IV methods within the Cox proportional hazard framework. Martinussen et al. (2017) introduced an IV-based approach for estimating structural cumulative survival models to capture time-varying effects on the survival function, and Kjaersgaard and Parner (2015) developed a pseudo-observation approach that accounts for the parametric contribution of covariates. In contrast, this work focuses on directly estimating HTE, which can be defined by different estimands, using nonparametric methods. This approach offers greater flexibility in capturing complex relationships between covariates and treatment effects and as we show can readily incorporate IV methods.
@@ 3 Problem Setup
for a known function $g(\cdot)$. For instance, with $g(\widetilde{T}_{i})=\min(\widetilde{T}_{i},h)$, $\tau(\cdot)$ represents the differences in restricted mean survival time (RMST), and with $g(\widetilde{T}_{i})=I(\widetilde{T}_{i}\geq h)$ it represents the differences in survival functions.
@@
are random forest-based weights, $\mathcal{N}_{b}(x)$ is the terminal node that contain $x$ in the $b$th tree, $B$ is the number of trees, and ||$\cdot$|| ⋅ | denotes the cardinality Athey et al. (2019).
@@ 4 The MISTR Procedure
In principle, Equation (2) above provides a path for flexible estimation of causal estimands. However, for time-to-event outcomes subject to right censoring, where $\widetilde{T}_{i}$ is only observed for a portion of the training sample, Eq. (2) is inapplicable. We propose an approach that relies on efficient imputation of event times for right-censored observations and thus allows us to make us of this equation. The core of MISTR includes the following steps: initially, RIST is employed to impute event times for censored observations. These imputed event times are then merged with the observed events to form a complete, uncensored dataset. Subsequently, $\tau(x)$ is estimated using Eq. (2). This procedure is replicated $A$ times, each time with different imputed event times, producing $A$ estimates, $\widehat{\tau}_{1}(x),\ldots,\widehat{\tau}_{A}(x)$. Our proposed estimate is the average of these $A$ estimates.
@@
21:$T_{i,a}\wedge h\leftarrow\min(T_{i,a},h)$, $a=1,\ldots,A,i=1,\ldots,n$.
22:$\delta_{i,a}^{h}\leftarrow I(T_{i,a}\leq h)$, $a=1,\ldots,A,i=1,\ldots,n$.
@@ 6.1 Benchmark Cases
We start by comparing the performance of MISTR to existing baselines. Cui et al. (2023) showed that their method is superior to the random survival forest (ishwaran_Rpackage_2019), S-learner (Künzel et al., 2019), enriched random survival forest (lu_estimating_2018), and the IPCW causal forest. Therefore, we focus on comparing our method with the CSF method of Cui et al. (2023). A comparison to additional baselines such as T-learner (Bo et al., 2024) is available in Table S3. First, we utilize Cui et al. (2023)’s benchmark (Settings 1–4). Then, we present new scenarios that go beyond the censoring probability assumptions required by CSF and IPCW approaches. These scenarios are characterized by censoring within a specific time interval (Settings 5, 10) higher censoring rates (Settings 6, 8, and 9), and unknown censoring mechanism (Setting 7).
For each scenario in settings 1–10, we generate a training set and two test sets: Random and Quantiles. The training set comprises $n_{train}=5000$ samples, $\{X_{i},W_{i},T_{i},\delta_{i}\}$, with $i=1,\ldots,n_{train}$. Each observation includes $p=5$ independent covariates generated from a U[0,1] distribution. In Setting 7, we sample two additional covariates from the same distribution, which are used only for sampling the censoring time $C_{i}$, and are unobserved during training or testing. The Random test set is generated in the same way as the training set with $n_{test}=5000$. The Quantiles test set comprises 21 observations, $\{X_{i},W_{i},T_{i},\delta_{i}\}$, where $i=1,\ldots,21$. Each $X_{i}$ includes $p=5$ covariates, all equal to a specific quantile value ${0.00,0.05,\ldots,0.95,1}$. For example, for $i=2$, $x_{2j}=0.1$ for $j=1,\ldots,5$. In Setting 7, two additional covariates, with values equal to the others, are included for sampling the censoring time $C_{i}$, e.g., for $i=2$, $x_{2j}=0.1$ for $j=1,\ldots,7$. True HTE values for all sets are approximated as the mean of 20,000 Monte Carlo samples. The sampling distributions of $C_{i}$ and $\widetilde{T}_{i}$ are detailed in Table S1, and the sampled populations are illustrated in Figure S1. The maximum time $t_{max}$, the horizon $h$, and the censoring rate are shown in Table S2.
@@ 6.2 Semi-simulated MIMIC Setting
The treatment assignment ($W$) is sampled from a Bernoulli$(0.5)$ distribution. Additionally, we use the covariate $X_{(36)}$ (standardized age), together with the first five lab test results (normal or abnormal binary indicators) $X_{(1)}-X_{(5)}$ to simulate failure time. The failure time is sampled from a Poisson distribution with $\lambda_{f}$ such that $\lambda_{f}=30+0.75(1-W)\bigg{[}\sum_{k=1}^{5}X_{(k)}+0.75X_{(36)}\bigg{]}-0.45W.$ Note that $\lambda_{f}$ is only influenced by part of the available baseline covariates, and that the HTE involves a combination of binary baseline covariates and a continuous covariate. Censoring time is sampled from a Poisson distribution with $\lambda_{c}={21,23,24.7,26.5,29}$ for Settings MIMIC 1–5, respectively, where the highest censoring rate is in MIMIC 1.
We randomly split the full dataset into 5 folds, numbered 1–5, each with $\sim 5030$ patients. The maximum time for the RIST is $t_{max}=29$, and the RMST horizon is $h=28$. The sampled populations are shown in Figure S4. We compare the HTE results of MISTR with those given by CSF (Cui et al., 2023) based on 5 replications, such that each fold is used either for training or testing. The mean and SE of the MSE across the 5 replications of the analysis are presented in Table 2. A comparison of the estimated effect versus the true effect for one test set is shown in Figure S5. Evidently, as observed in earlier experiments, at lower censoring percentages (e.g., MIMIC-5), both CSF and MISTR exhibit similar performance. However, as censoring increases, MISTR outperforms CSF. We use this setting to demonstrate the sensitivity of the variance to the bag size $\ell$ and to the number of imputations (Appendix B). Based on the results of these analyses, we recommend using $A\geq 100$. For the variance, a U-shaped relationship is expected due to the trade-off between the “within-imputation” and “between-imputation” components. Consequently, we recommend selecting the bag size using cross-validation.
@@ 6.3 Instrumental Variable Setting
In this sub-section we demonstrate the use and applicability of MISTR-IV. For each scenario, we generate a training set and a Random test set each comprises $n_{train}=n_{test}=5000$ observations, $\{X_{i},W_{i},Z_{i},U_{i},T_{i},\delta_{i}\}$, with $i=1,\ldots,n_{train}$ and $i=1,\ldots,n_{test}$, respectively. Each observation $i$ comprises $p=3$ independent baseline covariates $X_{i}$, and one hidden confounder $U_{i}$, both generated from a uniform distribution over $[0,1]$. The hidden confounder $U_{i}$ is used solely during data generation and remains unobserved during training and testing. The instrumental variable $Z_{i}$ is sampled from a Bernoulli$(0.5)$ distribution. For the treatment assignment, for each sample $i$ we first sample a variable $W_{i}^{(*)}$ from the distributions defined for each setting given the baseline covariates. The treatment assignment is then $W_{i}=1$ if $W_{i}^{(*)}>0.5$, and $W_{i}=0$ otherwise. We sample $\widetilde{T}_{i}$ and $C_{i}$ from Poisson distributions with parameters $\lambda_{\widetilde{T}_{i}}$ and $\lambda_{C_{i}}$ respectively, see Table S7 for details.
@@ Appendix C Additional Results - HIV Dataset Analysis
The HTEs for HIV-1 treatment options were in the range of $[0.71,4.99]$ months and $[0.59,4.95]$ months as estimated by MISTR and CSF, respectively. This means the ZVD+ddI is preferred over ZVD alone for all patients (no negative HTEs), albeit with varying benefits across patients. The Pearson correlation between the mean HTE of MISTR and CSF for HIV-1 was 0.995, i.e., both methods give similar HTE estimates, with smaller SE for MISTR, and with similar mean estimated SE. Similarly, for HIV-2, the HTEs were in range of $[1.00,4.30]$ months and $[0.91,4.22]$ months for MISTR and CSF, respectively with correlation of 0.997; and for HIV-3 in range of $[0.17,3.89]$ months and $[0.17,3.85]$ months for MISTR and CSF, respectively, with correlation of 0.996.
@@ C Poisson $\lambda_{C}=3$
Previous admission this month, n (%) No 23128 (91.93) 4599 (91.41) 4625 (91.93) 4653 (92.49) 4604 (91.51) 4647 (92.29)
@@
Marital status, n (%) Divorced 2041 (8.11) 409 (8.13) 429 (8.53) 418 (8.31) 398 (7.91) 387 (7.69)
@@
Red blood cells, n (%) Abnormal 19161 (76.16) 3843 (76.39) 3837 (76.27) 3856 (76.64) 3837 (76.27) 3788 (75.23)
@@
White blood cells, n (%) Abnormal 10003 (39.76) 2031 (40.37) 1987 (39.50) 1994 (39.63) 1945 (38.66) 2046 (40.64)
@@
Patient ID Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE
@@
Patient ID Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE Mean HTE Mean Est SE SE
@@
IV group, n (%) Control 3932 (32.6) 3932 (100.0) - -
Experiment 8125 (67.4) - 3953 (100.0) 4172 (100.0)

### v3

@@ 1 Introduction
In this paper, we introduce MISTR (Multiple Imputations for Survival Treatment Response), a new estimator for HTE estimation in right-censored survival data. Inspired by Recursively Imputed Survival Trees (RIST) (Zhu and Kosorok, 2012), MISTR eliminates the need to estimate the censoring distribution. This in turn allows the flexible use of methods which otherwise would not be applicable to right-censored data. We build upon this flexibility and show how one can create methods which are more robust, more accurate, and applicable to a broader settings compared with existing methods. Figure 1 presents our approach schematically, where censored observations are efficiently imputed multiple times, treatment effects are estimated for each imputed dataset, and the resulting estimators are then combined.
@@
Cui et al. (2023)
simulated settings
1–4 $\checkmark$ $\checkmark$ + $\checkmark$ +
Censoring probability
reaches low values
6 $\times$ $\times$ $\checkmark$ +
Censoring depends on
unobserved covariates
7 $\checkmark$ $\checkmark$ $\checkmark$ +
High censoring rate
8–9 $\checkmark$ $\checkmark$ $\checkmark$ +
Unobserved confounding
with IV adjustment
201–204 $\checkmark$ $\times$ $\checkmark$ +
@@ 2 Related Work
To estimate causal effects in the presence of confounding, IV methods can be employed (Angrist et al., 1996; Abadie, 2003; Hernán and Robins, 2006). An IV variable is associated with treatment assignment, affects the outcome only through the treatment, and is independent of unmeasured confounders. In survival data, IV methods have been widely studied. For example, Tchetgen Tchetgen et al. (2015) proposed a regression-based IV approach using additive hazard models, while Sørensen et al. (2019) and MacKenzie et al. (2014) applied IV methods within the Cox proportional hazard framework. Martinussen et al. (2017) introduced an IV-based approach for estimating structural cumulative survival models to capture time-varying effects on the survival function, and Kjaersgaard and Parner (2015) developed a pseudo-observation approach that accounts for the parametric contribution of covariates. In contrast, this work focuses on directly estimating HTE, which can be defined by different estimands, using nonparametric methods. This approach offers greater flexibility in capturing complex relationships between covariates and treatment effects and as we show can readily incorporate IV methods.
@@ 3 Problem Setup
for a known function $g(\cdot)$. For instance, with $g(\widetilde{T}_{i})=\min(\widetilde{T}_{i},h)$, where $h$ is a constant, $\tau(\cdot)$ represents the differences in restricted mean survival time (RMST), and with $g(\widetilde{T}_{i})=I(\widetilde{T}_{i}\geq h)$ it represents the differences in survival functions.
@@
are random forest-based weights, $\mathcal{N}_{b}(x)$ is the terminal node that contain $x$ in the $b$th tree, $B$ is the number of trees, and ||$\cdot$|| ⋅ | denotes the cardinality (Athey et al., 2019).
@@ 4 The MISTR Procedure
In principle, Equation (2) above provides a path for flexible estimation of causal estimands. However, for time-to-event outcomes subject to right censoring, where $\widetilde{T}_{i}$ is only observed for a portion of the training sample, Eq. (2) is inapplicable. We propose an approach that relies on efficient imputation of event times for right-censored observations and thus allows us to make use of this equation. The core of MISTR includes the following steps: initially, RIST is employed to impute event times for censored observations. These imputed event times are then merged with the observed events to form a complete, uncensored dataset. Subsequently, $\tau(x)$ is estimated using Eq. (2). This procedure is replicated $A$ times, each time with different imputed event times, producing $A$ estimates, $\widehat{\tau}_{1}(x),\ldots,\widehat{\tau}_{A}(x)$. Our proposed estimate is the average of these $A$ estimates.
@@
[…]
@@ 6.1 Benchmark Cases
We start by comparing the performance of MISTR to existing baselines. Cui et al. (2023) showed that their method is superior to the random survival forest (Ishwaran and Kogalur, 2019), S-learner (Künzel et al., 2019), enriched random survival forest (Lu et al., 2018), and the IPCW causal forest. Therefore, we focus on comparing our method with the CSF method of Cui et al. (2023). A comparison to additional baselines such as T-learner (Bo et al., 2024) is available in Table S3. First, we utilize Cui et al. (2023)’s benchmark (Settings 1–4). Then, we present new scenarios that go beyond the censoring probability assumptions required by CSF and IPCW approaches. These scenarios are characterized by censoring within a specific time interval (Settings 5, 10) higher censoring rates (Settings 6, 8, and 9), and unknown censoring mechanism (Setting 7).
[…]
@@ 6.2 Semi-simulated MIMIC Setting
The treatment assignment ($W$) is sampled from a Bernoulli(0.5)(0.5)( 0.5 ) distribution. Additionally, we use the covariate $X_{(36)}$ (standardized age), together with the first five lab test results (normal or abnormal binary indicators) $X_{(1)}-X_{(5)}$ to simulate failure time. The failure time is sampled from a Poisson distribution with $\lambda_{f}$ such that $\lambda_{f}=30+0.75(1-W)\bigg{[}\sum_{k=1}^{5}X_{(k)}+0.75X_{(36)}\bigg{]}-0.45W.$ Note that $\lambda_{f}$ is only influenced by part of the available baseline covariates, and that the HTE involves a combination of binary baseline covariates and a continuous covariate. Censoring time is sampled from a Poisson distribution with $\lambda_{c}={21,23,24.7,26.5,29}$ for Settings MIMIC 1–5, respectively, where the highest censoring rate is in MIMIC 1.
We randomly split the full dataset into 5 folds, numbered 1–5, each with 5030$\sim$ 5030∼ 5030 patients. The maximum time for the RIST is $t_{max}=29$, and the RMST horizon is $h=28$. The sampled populations are shown in Figure S4. We compare the HTE results of MISTR with those given by CSF (Cui et al., 2023) based on 5 replications, such that each fold is used either for training or testing. The mean and SE of the MSE across the 5 replications of the analysis are presented in Table 2. A comparison of the estimated effect versus the true effect for one test set is shown in Figure S5. Evidently, as observed in earlier experiments, at lower censoring percentages (e.g., MIMIC-5), both CSF and MISTR exhibit similar performance. However, as censoring increases, MISTR outperforms CSF. We use this setting to demonstrate the sensitivity of the variance to the bag size $\ell$ and to the number of imputations (Appendix B). Based on the results of these analyses, we recommend using $A\geq 100$. For the variance, a U-shaped relationship is expected due to the trade-off between the “within-imputation” and “between-imputation” components. Consequently, we recommend selecting the bag size using cross-validation.
@@ 6.3 Instrumental Variable Setting
In this sub-section we demonstrate the use and applicability of MISTR-IV. For each scenario, we generate a training set and a Random test set each comprises $n_{train}=n_{test}=5000$ observations, $\{X_{i},W_{i},Z_{i},U_{i},T_{i},\delta_{i}\}$, with $i=1,\ldots,n_{train}$ and $i=1,\ldots,n_{test}$, respectively. Each observation $i$ comprises $p=3$ independent baseline covariates $X_{i}$, and one hidden confounder $U_{i}$, both generated from a uniform distribution over [0,1][0,1][ 0 , 1 ]. The hidden confounder $U_{i}$ is used solely during data generation and remains unobserved during training and testing. The instrumental variable $Z_{i}$ is sampled from a Bernoulli(0.5)(0.5)( 0.5 ) distribution. For the treatment assignment, for each sample $i$ we first sample a variable $W_{i}^{(*)}$ from the distributions defined for each setting given the baseline covariates. The treatment assignment is then $W_{i}=1$ if $W_{i}^{(*)}>0.5$, and $W_{i}=0$ otherwise. We sample $\widetilde{T}_{i}$ and $C_{i}$ from Poisson distributions with parameters $\lambda_{\widetilde{T}_{i}}$ and $\lambda_{C_{i}}$ respectively, see Table S7 for details.
@@ Appendix C Additional Results - HIV Dataset Analysis
The HTEs for HIV-1 treatment options were in the range of [0.71,4.99][0.71,4.99][ 0.71 , 4.99 ] months and [0.59,4.95][0.59,4.95][ 0.59 , 4.95 ] months as estimated by MISTR and CSF, respectively. This means the ZVD+ddI is preferred over ZVD alone for all patients (no negative HTEs), albeit with varying benefits across patients. The Pearson correlation between the mean HTE of MISTR and CSF for HIV-1 was 0.995, i.e., both methods give similar HTE estimates, with smaller SE for MISTR, and with similar mean estimated SE. Similarly, for HIV-2, the HTEs were in range of [1.00,4.30][1.00,4.30][ 1.00 , 4.30 ] months and [0.91,4.22][0.91,4.22][ 0.91 , 4.22 ] months for MISTR and CSF, respectively with correlation of 0.997; and for HIV-3 in range of [0.17,3.89][0.17,3.89][ 0.17 , 3.89 ] months and [0.17,3.85][0.17,3.85][ 0.17 , 3.85 ] months for MISTR and CSF, respectively, with correlation of 0.996.
@@ C Poisson $\lambda_{C}=3$
Previous admission
this month, n (%)
No 23128 (91.93) 4599 (91.41) 4625 (91.93) 4653 (92.49) 4604 (91.51) 4647 (92.29)
@@
Marital
status, n (%)
Divorced 2041 (8.11) 409 (8.13) 429 (8.53) 418 (8.31) 398 (7.91) 387 (7.69)
@@
Red blood
cells, n (%)
Abnormal 19161 (76.16) 3843 (76.39) 3837 (76.27) 3856 (76.64) 3837 (76.27) 3788 (75.23)
@@
White blood
cells, n (%)
Abnormal 10003 (39.76) 2031 (40.37) 1987 (39.50) 1994 (39.63) 1945 (38.66) 2046 (40.64)
@@
Patient
ID
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
@@
Patient
ID
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
Mean
HTE
Mean
Est SE
SE
@@
IV group, n (%) Control 3932 (32.6) 3932 (100.0)
Experiment 8125 (67.4)
3953 (100.0) 4172 (100.0)